# Fluctuations in random field Ising models

### v1

@@ 2.4 Random field Ising models
Note that the above corollary applies to any vector $\mathbf{q}$ which is delocalized, in the sense that $\|\mathbf{q}\|_{\infty}\to 0$. We stress that without this condition CLT fails even when $\operatorname{\mathnormal{\bm{\sigma}}}$ is a vector of i.i.d. random variables. Thus, essentially there is no restriction on the vector ${\bf q}$. We can also allow for dependence across the entries of the rows of the data matrix ${\bf Z}:=(Z_{k,i})_{k\in[N],i\in[n]}$, at the expense of demanding that $\mathbf{q}$ is an approximate eigenvector of the corresponding population covariance matrix (after removing the diagonal entries).
@@
We have presented Theorems 2.3, 2.4, and 2.5, that allow general $A_{n}$, $\mathbf{c}$, $\mu_{i}$s. In particular, these results do not rely on the assumption that $c_{1},\ldots,c_{n}$ are i.i.d. Thus, we expect them to be applicable to more general quadratic interaction models. A particularly challenging case is when $\operatorname{\mathbb{P}}$ is the posterior distribution in a linear regression model with Gaussian errors. In that case, the corresponding field vector is neither identically distributed nor does it possess independent coordinates, and the base measures $\mu_{i}$ are heterogeneous. We plan to pursue this problem in the future.
@@ 2.5 Discussion
2.6 Organization of proofs
We prove the main Theorems 2.3, 2.4, and 2.5 in Section 3. The proof of 2.3 is based on a recursive computation that crucially utilizes the high-temperature condition MHT (see (2.2)). The proof 2.4 is based on Stein’s method of exchangeable pairs and a tight control of the local fields (1.5) using contractions in the $(4,4)$-operator norm. For these arguments, we utilize the approximate eigenpair assumption for $(\mathbf{A}_{n},\mathbf{q})$ as well as the MHT condition. 2.5 follows from doing a Taylor expansion of the Mean-Field optimizers $\mathbf{u}$ and simplifying the expression for $\mathbf{q}^{\top}\mathbf{u}$ using the approximate eigenpair condition.
In Section 4, we prove the theoretical claims for our results on random field Ising models. Section 5 is devoted to the proofs of technical lemmas that were previously used to prove our main results. Finally, we prove some technical matrix lemmas in the Appendix Section 6.
@@ 3 Proof of main results
$\mathbf{v}^{\top}\nabla^{2}F(\operatorname{\mathbf{z}})\mathbf{v}$ $=\left(\mathbf{v}^{\top}\mathbf{A}_{n}\mathbf{v}-\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\frac{v_{i}^{2}}{\psi_{i}^{\prime\prime}(\phi_{i}^{-1}(z_{i}))}\right)$
@@
where the final inequality follows from the WHT condition (2.1). Therefore $F(\cdot)$ is strictly concave, and so, there exists a unique maximizer $\mathbf{u}$ in the convex domain $[-1,1]^{n}$. This maximizer cannot be at the boundary of $[-1,1]^{n}$, since
$\frac{\partial}{\partial z_{i}}F(\operatorname{\mathbf{z}})=\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)z_{j}-\phi_{i}(z_{i})$
@@
For a bounded measurable function $g(\cdot)$ on $[-1,1]$, let $\mathbf{b}^{(g)}:=(b^{(g)}_{1},\cdots,b^{(g)}_{n})\in[-1,1]^{n}$ denote the vector of conditional expectations of $g(\cdot)$ under model (1.1), defined by
@@
Suppose $\mathbf{A}_{n}$ satisfies $\lVert\mathbf{A}_{n}\rVert\leq h<\infty$. Let $\bm{\gamma}:=(\gamma_{1},\ldots,\gamma_{n})\in\mathbb{R}^{n}$ be a deterministic $p$-dimensional vector. Then, for any bounded measurable function $g(\cdot)$ on $[-1,1]$, the following inequalities hold:
@@
    $\operatorname{\mathbb{P}}\left(\bigg{|}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\gamma_{i}\big{(}g(\sigma_{i})-b^{(g)}_{i}\big{)}\bigg{|}>t\right)\leq 2\exp\left(-\frac{at^{2}}{\|\gamma\|^{2}}\right)$
    for some constant $a>0$ depending only on $h$. Consequently, for $r>0$, we have
@@
    The hidden constant here depends on $h$ but not on $g(\cdot)$ and $\bm{\gamma}$.
While similar results have been established in the literature (see for e.g. [19, Lemma 2.1] and [4, Lemma 2.10]), the key distinguishing features of Lemma 3.1 are that we impose no restrictions on the external field $\mathbf{c}$ and the weight vector $\bm{\gamma}$.
@@
$|T_{(\ell)}|\leq\sqrt{n}\|\mathbf{C}_{n}\|^{\ell}\|{\mathbf{q}}\|\leq\sqrt{n}\|\mathbf{A}_{n}\|^{\ell}\leq\sqrt{n}\rho^{\ell}\stackrel{{\scriptstyle\ell\to\infty}}{{\to}}0.$ (3.3)
@@
$\leq\sum_{\ell=0}^{\infty}\bigg{|}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{(\ell)}(\sigma_{i}-u_{i})-\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{(\ell+1)}(\sigma_{i}-u_{i})\bigg{|}$
$\leq\sum_{\ell=0}^{\infty}\bigg{|}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{(\ell)}(\sigma_{i}-b_{i})\bigg{|}+\sum_{\ell=0}^{\infty}\bigg{|}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{(\ell)}(m_{i}-s_{i})^{2}\psi_{i}^{\prime\prime\prime}(\xi_{i}+s_{i})\bigg{|}.$
@@
$\leq\sum_{\ell=0}^{\infty}\operatorname{\mathbb{P}}\left(\bigg{|}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{(\ell)}(\sigma_{i}-b_{i})\bigg{|}\geq ta_{\ell}\right)+\sum_{\ell=0}^{\infty}\operatorname{\mathbb{P}}\left(8\lVert\mathbf{m}-\mathbf{s}\rVert_{4}^{2}\geq ct(\ell+1)\right).$ (3.4)
@@
$\sum_{\ell=0}^{\infty}\operatorname{\mathbb{P}}\left(\bigg{|}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{(\ell)}(\sigma_{i}-b_{i})\bigg{|}\geq ta_{l}\right)$ $\leq 2\sum_{\ell=0}^{\infty}\exp\left(-c\frac{t^{2}a_{\ell}^{2}}{\rho^{2\ell}}\right)=2\sum_{\ell=0}^{\infty}\exp\left(-ct^{2}(\ell+1)^{2}\right)$
$\leq 2\frac{\exp(-ct^{2})}{1-\exp(-c)},$ (3.5)
where the last inequality holds for $t\geq 1$. We now split the proof in two parts.
@@
[…]
The conclusion now follows by invoking Lemma 3.2(b).
@@
$\leq\sum_{\ell=0}^{\infty}\exp(-\vartheta_{n}ct(\ell+1))\left(1+\sum_{r=1}^{\infty}n^{\frac{r}{2}-1}\frac{(8\vartheta_{n})^{r}}{r!}\operatorname{\mathbb{E}}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}(m_{i}-s_{i})^{2r}\right)$
@@
The first display above follows from Markov’s inequality. The second display uses the power series expansion of the exponential function. The third display is a consequence of Hölder’s inequality. For the fourth display, we lever Lemma 3.2(c). The fifth display is a product of the standard lower bound $r!\geq\sqrt{2\pi}r^{r+1/2}e^{-r}$. The final inequality is immediate from the choice of $\vartheta_{r}$, and the assumption $t\geq\sqrt{n}\operatorname{\alpha_{\mathnormal{n}}}$. The above display, combined with (3) and (3), concludes the proof. ∎
@@
For all $i=1,\ldots,n$, let $\nu_{i}=\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)\psi_{j}^{\prime}(c_{j})$ be as in 2.4, and $\bm{\nu}=(\nu_{1},\ldots,\nu_{n})$. Recall the definition of ${\bm{s}}$ from Lemma 2.2.
@@
Note that $|\Delta_{n}|\leq 2\lVert\mathbf{q}\rVert_{\infty}$, and recall the definition of $W_{n}$ from 2.4. By combining (3.6) with [35, Corollary 2.1], we have
$d_{KS}(T_{n}^{*}(\operatorname{\mathnormal{\bm{\sigma}}}),W_{n})\lesssim\operatorname{\mathbb{E}}\left|1-\frac{n}{2\upsilon_{n}}\operatorname{\mathbb{E}}(\Delta_{n}^{2}\mid T_{n}^{*}(\operatorname{\mathnormal{\bm{\sigma}}}))\right|+n\operatorname{\mathbb{E}}\left[\sum_{a=1}^{3}|H_{a}|\right]+\lVert\mathbf{q}\rVert_{\infty}.$ (3.7)
@@
$\frac{p}{2}\operatorname{\mathbb{E}}(\Delta_{n}^{2}\mid\operatorname{\mathnormal{\bm{\sigma}}})$ $=\frac{1}{2}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{2}\operatorname{\mathbb{E}}\left(\sigma_{i}^{2}+(\sigma_{i}^{\prime})^{2}-2\sigma_{i}\sigma^{\prime}_{i}\mid\operatorname{\mathnormal{\bm{\sigma}}}\right)$
@@
$\operatorname{\mathbb{E}}\bigg{|}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{2}\left(\sigma_{i}^{2}-\operatorname{\mathbb{E}}[\sigma_{i}^{2}|\sigma_{j},\ j\neq i]\right)\bigg{|}+\operatorname{\mathbb{E}}\bigg{|}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{2}\operatorname{\mathbb{E}}[\left(\sigma_{i}-b_{i}\right)b_{i}]\bigg{|}\lesssim\sqrt{\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{4}}\leq\lVert\mathbf{q}\rVert_{\infty}.$
@@
By combining the relevant bounds from above, we get:
$\operatorname{\mathbb{E}}\left|1-\frac{n}{2\upsilon_{n}}\operatorname{\mathbb{E}}(\Delta_{n}^{2}\mid T_{n}^{*}(\operatorname{\mathnormal{\bm{\sigma}}}))\right|\lesssim\frac{1}{\upsilon_{n}}\left(\lVert\mathbf{q}\rVert_{\infty}+\sqrt{n\alpha_{n}^{2}}+\sqrt{R_{3n}}\right).$
@@
where the last inequality follows from 2.3. In order to bound $\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\gamma_{i}^{2}$, we note that for some $\{\xi_{j}\}_{1\leq i\leq n}$,
@@
$\leq 2\bigg{|}\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)q_{j}(\psi_{j}^{\prime\prime}(c_{j})-\upsilon_{n})\bigg{|}^{2}+16\left(\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)^{2}q_{j}^{2}\right)\left(\sum_{\mathnormal{j}=1}^{\mathnormal{n}}s_{j}^{2}\right)$
and therefore, by applying Lemma 3.3, we get:
$\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\gamma_{i}^{2}$ $\lesssim\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\bigg{|}{\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)q_{j}(\psi_{j}^{\prime\prime}(c_{j})-\upsilon_{n})}\bigg{|}^{2}+\left(\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)^{2}q_{j}^{2}\right)\left(\sum_{\mathnormal{j}=1}^{\mathnormal{n}}t_{j}^{2}\right)$
@@
Finally, let us bound $H_{3}$. Recall that $\bm{\epsilon}=\mathbf{A}_{n}\mathbf{q}-\lambda_{n}\mathbf{q}$. We again use 2.3 to see
@@
$d_{KS}(aX+b,\operatorname{N}(0,\tau))\lesssim_{K}|a^{2}-1|+|b|+d_{KS}(X,\operatorname{N}(0,\tau)).$ (3.9)
@@
    where $e_{i}:=\frac{s_{i}^{2}}{2}\psi_{i}^{\prime\prime\prime}(\xi_{i}+c_{i})$, for some reals $\{\xi_{i}\}_{1\leq i\leq n}$. Then, with $\Psi^{\prime}(\mathbf{c})=(\psi_{1}^{\prime}(c_{1}),\ldots,\psi_{n}^{\prime}(c_{n}))$, we have
@@
    Since $01-\lambda_{n}\upsilon_{n}\geq 1-\rho>$ 0, it suffices to bound the terms in the RHS of (3.10). Towards this direction, note that the first term can be controlled by writing
@@
    $\mathbf{q}^{\top}\operatorname{\mathbf{e}}\,\leq\|\operatorname{\mathbf{e}}\|\lesssim\sqrt{\sum_{\mathnormal{i}=1}^{\mathnormal{n}}s_{i}^{4}}\,\lesssim\sqrt{\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\nu_{i}^{4}}\,=\sqrt{R_{3n}}\,.$
@@
    Proceeding to verify part (b), invoking Lemma 3.4 with the choices
    $X=\mathbf{q}^{\top}(\bm{\sigma}-\mathbf{u}),\quad\tau=\frac{1}{1-\lambda\upsilon_{n}}\in[1-\rho,1/(1-\rho)],\quad a=1,\quad b=\mathbf{q}^{\top}\left(\mathbf{u}-\frac{\Psi^{\prime}(\mathbf{c})}{1-\lambda_{n}\upsilon_{n}}\right)$

### v2

@@ 2.5 Random field Ising models
The above corollary applies to any vector $\mathbf{q}$ which is delocalized, in the sense that $\|\mathbf{q}\|_{\infty}\to 0$. We stress that without this condition CLT fails even when $\operatorname{\mathnormal{\bm{\sigma}}}$ is a vector of i.i.d. random variables. Thus, essentially there is no restriction on the vector $\bf q$}. We can also allow for dependence across the entries of the rows of the data matrix ${\bf Z}:=(Z_{k,i})_{k\in[N],i\in[n]}$, at the expense of demanding that $\mathbf{q}$ is an approximate eigenvector of the corresponding population covariance matrix (after removing the diagonal entries).
Note that we work under an asymptotic scaling where the matrix $\bf Z$} has significantly more rows than columns, as opposed to the proportional asymptotics with $n/N\to\gamma\in(0,\infty).$ Also, note that we can allow an arbitrary diluting of our Hopfield model. Indeed, this is because we do not re-scale $\mathbf{A}_{n}$ by the average degree of the graph $\mathbf{G}_{n}$, in contrast to what is done in [9], which considers the randomly diluted Hopfield model where $\mathbf{G}_{n}$ is the Erdős–Rényi random graph with connection probability $p$, and divides each $\mathbf{A}_{n}(i,j)$ by $p$. Our main results directly apply to this version of the diluted Hopfield model from [9], as long as the Erdős–Rényi probability $p$ is bounded away from 0 (and more generally to dense graphs). It remains unclear whether our results apply to the diluted model for non-dense graphs, scaled by the average degree. The main challenge in this direction is to verify the MHT condition.
@@
[…]
@@ 2.6 Discussion
Organization of proofs  We prove the main Theorems 2.3, 2.4, and 2.5 in Section 3. In Section 4, we prove the theoretical claims for our results on random field Ising models. Section 5 is devoted to the proofs of technical lemmas that were previously used to prove our main results. Finally, we prove some technical matrix lemmas in the Appendix Section 6.
@@ 3 Proof of main results
$\mathbf{v}^{\top}\nabla^{2}F(\operatorname{\mathbf{z}})\mathbf{v}$ $=\left(\mathbf{v}^{\top}\mathbf{A}_{n}\mathbf{v}-\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\frac{v_{i}^{2}}{\psi_{i}^{\prime\prime}(\phi_{i}(z_{i}))}\right)$
@@
where the final inequality follows from the WHT condition (2.1). Therefore $F(\cdot)$ is strictly concave, and so, there exists a unique maximizer $\mathbf{u}$ in the convex domain $1,1]^{n}$. This maximizer cannot be at the boundary of $1,1]^{n}$, since
$\frac{\partial}{\partial z_{i}}F(\operatorname{\mathbf{z}})=\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)z_{j}+c_{i}-\phi_{i}(z_{i})$
@@
For a bounded measurable function $g(\cdot)$ on [1,1][-1,1], let $\mathbf{b}^{(g)}:=(b^{(g)}_{1},\cdots,b^{(g)}_{n})^{\top}\in[-1,1]^{n}$ denote the vector of conditional expectations of $g(\cdot)$ under model (1.1), defined by
@@
Suppose $\mathbf{A}_{n}$ satisfies $\lVert\mathbf{A}_{n}\rVert\leq h<\infty$. Let $\bm{\gamma}:=(\gamma_{1},\ldots,\gamma_{n})^{\top}\in\mathbb{R}^{n}$ be a deterministic $n$-dimensional vector. Then, for any bounded measurable function $g(\cdot)$ on [1,1][-1,1], the following inequalities hold:
@@
    $\operatorname{\mathbb{P}}\left(\bigg|\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\gamma_{i}\big(g(\sigma_{i})-b^{(g)}_{i}\big)\bigg|>t\right)\leq 2\exp\left(-\frac{at^{2}}{\|\bm{\gamma}\|^{2}}\right)$
    for some constant $a>0$ depending only on $h$ and $\|g\|_{\infty}$. Consequently, for $r>0$, we have
@@
    The hidden constant here depends on $h$ and and $\|g\|_{\infty}$ but not on $\bm{\gamma}$.
While similar results have been established in the literature (see for e.g. [29, Lemma 2.1] and [4, Lemma 2.10]), the key distinguishing features of Lemma 3.1 are that we impose no restrictions on the external field $\mathbf{c}$ and the weight vector $\bm{\gamma}$.
@@
$|T_{(\ell)}|\leq 2\sqrt{n}\|\mathbf{C}_{n}\|^{\ell}\|{\mathbf{q}}\|\leq 2\sqrt{n}\|\mathbf{A}_{n}\|^{\ell}\leq 2\sqrt{n}\rho^{\ell}\stackrel{{\scriptstyle\ell\to\infty}}{{\to}}0.$ (3.3)
@@
$\leq\sum_{\ell=0}^{\infty}\bigg|\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{(\ell)}(\sigma_{i}-b_{i})+\frac{1}{2}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{(\ell+1)}(m_{i}-s_{i})^{2}\psi_{i}^{\prime\prime\prime}(\xi_{i}+c_{i})\bigg|$
$\leq\sum_{\ell=0}^{\infty}\bigg|\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{(\ell)}(\sigma_{i}-b_{i})\bigg|+\frac{1}{2}\sum_{\ell=0}^{\infty}\bigg|\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{(\ell)}(m_{i}-s_{i})^{2}\psi_{i}^{\prime\prime\prime}(\xi_{i}+s_{i})\bigg|.$
@@
$\leq\sum_{\ell=0}^{\infty}\operatorname{\mathbb{P}}\left(\bigg|\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{(\ell)}(\sigma_{i}-b_{i})\bigg|\geq ta_{\ell}\right)+\sum_{\ell=0}^{\infty}\operatorname{\mathbb{P}}\left(4\lVert\mathbf{m}-\mathbf{s}\rVert_{4}^{2}\geq ct(\ell+1)\right).$ (3.4)
@@
$\sum_{\ell=0}^{\infty}\operatorname{\mathbb{P}}\left(\bigg|\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{(\ell)}(\sigma_{i}-b_{i})\bigg|\geq ta_{l}\right)$ $\leq 2\sum_{\ell=0}^{\infty}\exp\left(-a\frac{t^{2}a_{\ell}^{2}}{\rho^{2\ell}}\right)$
$=2\sum_{\ell=0}^{\infty}\exp\left(-ac^{2}t^{2}(\ell+1)^{2}\right)$
$\leq 2\sum_{\ell=0}^{\infty}\exp\left(-ac^{2}t^{2}(\ell+1)\right)$
$\leq 2\frac{\exp(-ac^{2}t^{2})}{1-\exp(-ac^{2})},$ (3.5)
where the last inequality holds for $t\geq 1$, and $a>0$ is the constant from Lemma 3.1. We now split the proof in two parts.
@@
$\operatorname{\mathbb{E}}\big(T_{n}^{*}(\operatorname{\mathnormal{\bm{\sigma}}})\big)^{2}$ $=8\int_{0}^{\infty}t\,\operatorname{\mathbb{P}}\left(\bigg|\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}(\sigma_{i}-u_{i})\bigg|\geq 2t\right)\,dt$
$\leq 8\int_{0}^{1}tdt+16\int_{1}^{\infty}t\frac{\exp(-ac^{2}t^{2})}{1-\exp(-ac^{2})}dt+8\sum_{l=0}^{\infty}\int_{0}^{\infty}t\operatorname{\mathbb{P}}\Big(\frac{4\|\mathbf{m}-\mathbf{s}\|_{4}^{2}}{c(\ell+1)}\geq t\Big)dt$
$=4+\frac{16}{1-\exp(-ac^{2})}\int_{1}^{\infty}t\exp(-ac^{2}t^{2})\,dt+\frac{64}{c^{2}}\operatorname{\mathbb{E}}\|\mathbf{m}-\mathbf{s}\|_{4}^{4}\sum_{\ell=0}^{\infty}(\ell+1)^{-2}.$
Here, the first and final equality follow from the identity $\operatorname{\mathbb{E}}X_{n}^{2}=2\int_{0}^{\infty}t\operatorname{\mathbb{P}}(|X_{n}|\geq t)dt$. The conclusion now follows by invoking Lemma 3.2(b).
@@
$\leq\sum_{\ell=0}^{\infty}\exp(-\vartheta_{n}ct(\ell+1))\left(1+8\vartheta_{n}\sqrt{\operatorname{\mathbb{E}}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}(m_{i}-s_{i})^{4}}+\sum_{r=2}^{\infty}n^{\frac{r}{2}-1}\frac{(8\vartheta_{n})^{r}}{r!}\operatorname{\mathbb{E}}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}(m_{i}-s_{i})^{2r}\right)$
@@
The first display above follows from Markov’s inequality. The second display uses the power series expansion of the exponential function. The third display is a consequence of Hölder’s inequality (for the summands indexed by $r\geq 2$). For the fourth display, we use Lemma 3.2(c) to control both expectations. The fifth display is a product of the standard lower bound $r!\geq\sqrt{2\pi}r^{r+1/2}e^{-r}$. The final inequality is immediate from the choice of $\vartheta_{n}$, and the assumption $t\geq\sqrt{n}\operatorname{\alpha_{\mathnormal{n}}}$. The above display, combined with (3) and (3), concludes the proof. ∎
@@
For all $i=1,\ldots,n$, let $\nu_{i}=\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)\psi_{j}^{\prime}(c_{j})$ be as in 2.4, and $\bm{\nu}=(\nu_{1},\ldots,\nu_{n})^{\top}$. Recall the definition of $\bm{s}$} from Lemma 2.2.
@@
Note that $|\Delta_{n}|\leq 2\lVert\mathbf{q}\rVert_{\infty}$, and recall the definition of $W_{n}$ from 2.4.
Now we apply [45], which provides a KS distance bound between a random variable $S_{n}^{*}$ and a standard normal. Recall from the statement of 2.4 that the constants $\rho,\kappa>0$ satisfy $|\lambda_{n}|\leq\rho<1,\upsilon_{n}\geq\kappa$. For notational convenience, set $S_{n}^{*}(\mathnormal{\bm{\sigma}}):=T_{n}^{*}(\mathnormal{\bm{\sigma}})\sqrt{\frac{1-\lambda_{n}\upsilon_{n}}{\upsilon_{n}}}$ as a re-scaled version of $T_{n}^{*}$ and note that
$\operatorname{\mathbb{E}}(S_{n}^{*}(\mathnormal{\bm{\sigma}})-S_{n}^{*}(\mathnormal{\bm{\sigma}}^{\prime})\mid S_{n}^{*}(\mathnormal{\bm{\sigma}}))=g(S_{n}^{*}(\mathnormal{\bm{\sigma}}))-\sqrt{\frac{1-\lambda_{n}\upsilon_{n}}{\upsilon_{n}}}\operatorname{\mathbb{E}}(H_{1}+H_{2}+H_{3}\mid S_{n}^{*}(\mathnormal{\bm{\sigma}})).$
Set $\delta_{n}:=2\lVert\mathbf{q}\rVert_{\infty}\sqrt{\frac{1-\lambda_{n}\upsilon_{n}}{\upsilon_{n}}}$ so that $|S_{n}^{*}(\mathnormal{\bm{\sigma}})-S_{n}^{*}(\mathnormal{\bm{\sigma}}^{\prime})|\leq\delta_{n}.$ Noting that $\operatorname{\mathbb{E}}|S_{n}^{*}(\mathnormal{\bm{\sigma}})|=\operatorname{\mathbb{E}}|T_{n}^{*}(\mathnormal{\bm{\sigma}})|\sqrt{\frac{1-\lambda_{n}\upsilon_{n}}{\upsilon_{n}}}\lesssim 1$ by using the preliminary bound (3.6), $\upsilon_{n}\geq\kappa>0$, and $1-\lambda_{n}\upsilon_{n}\leq 1+\rho$, we can apply [45, Corollary 2.1] (in the second line below), which gives
$d_{KS}(T_{n}^{*}(\operatorname{\mathnormal{\bm{\sigma}}}),W_{n})$ $=d_{KS}(S_{n}^{*}(\mathnormal{\bm{\sigma}}),\operatorname{N}(0,1))$
$\leq\operatorname{\mathbb{E}}\left|1-\frac{n}{2(1-\lambda_{n}\upsilon_{n})}\operatorname{\mathbb{E}}((S_{n}^{*}(\mathnormal{\bm{\sigma}})-S_{n}^{*}(\mathnormal{\bm{\sigma}}^{\prime}))^{2}\mid S_{n}^{*}(\operatorname{\mathnormal{\bm{\sigma}}}))\right|+\frac{n\operatorname{\mathbb{E}}\left[\sum_{a=1}^{3}|H_{a}|\right]}{\sqrt{\upsilon_{n}(1-\lambda_{n}\upsilon_{n})}}+O(\delta_{n})$
$\lesssim\operatorname{\mathbb{E}}\left|1-\frac{n}{2\upsilon_{n}}\operatorname{\mathbb{E}}(\Delta_{n}^{2}\mid T_{n}^{*}(\operatorname{\mathnormal{\bm{\sigma}}}))\right|+{n\operatorname{\mathbb{E}}\left[\sum_{a=1}^{3}|H_{a}|\right]}+\lVert\mathbf{q}\rVert_{\infty}.$ (3.7)
For the final line, we again used $\upsilon_{n}\geq\kappa>0$ and $1-\lambda_{n}\upsilon_{n}\geq 1-\rho>0$ to simplify the denominator.
@@
$\frac{n}{2}\operatorname{\mathbb{E}}(\Delta_{n}^{2}\mid\operatorname{\mathnormal{\bm{\sigma}}})$ $=\frac{1}{2}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}^{2}\operatorname{\mathbb{E}}\left(\sigma_{i}^{2}+(\sigma_{i}^{\prime})^{2}-2\sigma_{i}\sigma^{\prime}_{i}\mid\operatorname{\mathnormal{\bm{\sigma}}}\right)$
@@
[…]
@@
By combining the relevant bounds from above and recalling that $\upsilon_{n}\geq\kappa$, we get:
$\operatorname{\mathbb{E}}\left|1-\frac{n}{2\upsilon_{n}}\operatorname{\mathbb{E}}(\Delta_{n}^{2}\mid T_{n}^{*}(\operatorname{\mathnormal{\bm{\sigma}}}))\right|\lesssim\left(\lVert\mathbf{q}\rVert_{\infty}+\sqrt{n\alpha_{n}^{2}}+\sqrt{R_{3n}}\right).$
@@
where the last inequality follows from 2.3. Note that we used $\max(1,n\alpha_{n}^{2})=1$ to simplify the RHS, which follows from the assumption $\sqrt{n}\operatorname{\alpha_{\mathnormal{n}}}\leq 1$ (see one line before (3.6)). In order to bound $\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\gamma_{i}^{2}$, we note that for some $\{\xi_{j}\}_{1\leq i\leq n}$,
@@
$\leq 2\bigg|\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)q_{j}(\psi_{j}^{\prime\prime}(c_{j})-\upsilon_{n})\bigg|^{2}+128\left(\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)^{2}q_{j}^{2}\right)\left(\sum_{\mathnormal{j}=1}^{\mathnormal{n}}s_{j}^{2}\right),$
where we use the bound $|\psi_{j}^{\prime\prime\prime}(\cdot)|\leq 8$. Therefore, by applying Lemma 3.3, we get:
$\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\gamma_{i}^{2}$ $\lesssim\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\bigg|{\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)q_{j}(\psi_{j}^{\prime\prime}(c_{j})-\upsilon_{n})}\bigg|^{2}+\left(\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)^{2}q_{j}^{2}\right)\left(\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\nu_{j}^{2}\right)$
@@
Finally, we bound $H_{3}$. Recall that $\bm{\epsilon}=\mathbf{A}_{n}\mathbf{q}-\lambda_{n}\mathbf{q}$. We again use 2.3 and $\max(1,n\alpha_{n}^{2})=1$ to see
@@
$d_{KS}(aX+b,\operatorname{N}(0,\tau))\lesssim|a^{2}-1|+|b|+d_{KS}(X,\operatorname{N}(0,\tau))$ (3.9)
where the hidden constant only depends on $K$.
@@
    where $e_{i}:=\frac{s_{i}^{2}}{2}\psi_{i}^{\prime\prime\prime}(\xi_{i}+c_{i})$, for some reals $\{\xi_{i}\}_{1\leq i\leq n}$. Then, with $\Psi^{\prime}(\mathbf{c})=(\psi_{1}^{\prime}(c_{1}),\ldots,\psi_{n}^{\prime}(c_{n}))^{\top}$, we have
@@
    Since $1-\lambda_{n}\upsilon_{n}\geq 1-\rho>0$, it suffices to bound the terms in the RHS of (3.10). Towards this direction, note that the first term can be controlled by writing
@@
    $|\mathbf{q}^{\top}\operatorname{\mathbf{e}}|\,\leq\|\operatorname{\mathbf{e}}\|\lesssim\sqrt{\sum_{\mathnormal{i}=1}^{\mathnormal{n}}s_{i}^{4}}\,\lesssim\sqrt{\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\nu_{i}^{4}}\,=\sqrt{R_{3n}}\,.$
@@
    Proceeding to verify part (b), recalling that $|\lambda_{n}|\leq\rho<1$ and $\upsilon_{n}\leq 1$, we have
    $1-\rho\leq 1-\lambda_{n}\upsilon_{n}\leq 1+\rho.$
    Also recall that we assume $\upsilon_{n}\geq\kappa$. Invoking Lemma 3.4 with the choices
    $X=\mathbf{q}^{\top}(\bm{\sigma}-\mathbf{u}),\quad\tau=\frac{\upsilon_{n}}{1-\lambda_{n}\upsilon_{n}}\in[\kappa(1-\rho),1+\rho],\quad a=1,\quad b=\mathbf{q}^{\top}\left(\mathbf{u}-\frac{\Psi^{\prime}(\mathbf{c})}{1-\lambda_{n}\upsilon_{n}}\right)$